# Strongly magnetized accretion with low angular momentum produces a weak jet

### v1

@@ 4 Conclusion and discussion
We have restricted our study to an initial magnetic field in the laboratory frame which is strictly poloidal and aligned with the spin axis of the BH. However, simulations with a Bondi radius of $200r_{g}$ and zero angular momentum (Ressler et al., 2021) have shown that tilting the magnetic field affects the resulting jet power. It remains unclear whether this carries over to the larger Bondi radius regime of most astrophysical interest. Therefore, in future studies, it would be interesting to explore the impact of tilting the magnetic field, as well as gas rotation, relative to the spin axis of the BH. These model problems could help in understanding the connection between the finite circularization radius problem studied here and simulations that begin with an equilibrium torus initial condition, in which tilting the magnetic field and plasma rotation axis relative to the BH spin axis can dramatically change the dynamics (e.g. Liska et al., 2018, 2019). Inflowing gas with finite circularization radius likely serves as a better initial condition for accretion of hot gas in the centers of galaxies than an equilibrium torus (e.g. Cuadra et al., 2006; Guo et al., 2024).
@@
The results presented here on the interplay between accretion and jets in GRMHD simulations of spherical accretion build on those of Ressler et al. (2021), Kwan et al. (2023), and Lalakos et al. (2024); we have extended these author’s results by varying the black hole spin, the initial gas angular momentum, the Bondi radius, and the initial plasma-$\beta$. For zero gas angular momentum, we find a similar dynamical evolution to that reported by Lalakos et al. (2024) – the initial emergence of a powerful jet with $\eta>100\%$, followed by its eventual disappearance as the normalized magnetic flux at the horizon ($\phi_{\rm BH}$) saturates below the typical MAD value. However, we attribute this behavior to the turbulence driven by eruption episodes, rather than the jet’s current-driven kink instability, as proposed by Lalakos et al. (2024). This interpretation is supported by our investigation of the $a=0$ case, where no jets are produced. Even in the absence of jets, the coherence of the accreted magnetic field is disturbed, leading to less efficient magnetic flux accumulation. In addition, if the transitions from $\eta>100\%$ to a low $\eta$ state were driven by jet feedback on large scales, we would expect the transition to take longer for a larger Bondi radius and for the accumulated magnetic flux to plausibly depend on the Bondi radius. Neither of these is true in our simulations. For example, our results show that the duration of the high-$\eta$ jet phase depends weakly on the Bondi radius, as seen in the initial spike in $\eta$ in Fig. 9. This is consistent with the disruption of magnetic flux accumulation originating in feedback from magnetic eruptions close to the black hole, rather than from larger radii. Once the strong jet is disrupted, however, the remaining weaker jets are indeed more prone to the kink instability, which generates weak jets that dissipate their energy over a large solid angle (Fig. 3).
@@
One possible application of our findings is to accretion onto Sgr A*. In their simulations of accretion onto Sgr A* from the surrounding stellar winds, Ressler et al. (2023) found periods of time with large dimensionless magnetic flux and strong jets, and periods of time with weak jets. These may be analogous to the high and low angular momentum states found here. More broadly, our results suggest that in the low-angular momentum state explored in this work, observations should indicate the presence of strong magnetic fields even in the absence of a powerful jet. This is potentially consistent with the EHT and GRAVITY evidence for strong magnetic fields in Sgr A* (Event Horizon Telescope Collaboration et al., 2024; GRAVITY Collaboration et al., 2018; Gravity Collaboration et al., 2023), but the lack of definitive evidence for a powerful jet. As a check on this interpretation, we performed ray-tracing calculations of the synchrotron emission in the strong and weak jet states found in this work (as described in § 3.1). We find that the linear and circular polarization levels are indeed similar in both cases. Low angular momentum in the inflowing gas could also help explain the surprising lack of azimuthal asymmetry in the EHT image of Sgr A* (Cole Faggert et al., 2024).
AG and AP thank Misha Medvedev for fruitful discussions; we also thank Sasha Tchekhovskoy for useful conversations. The simulations presented in this work were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center and Visualization Laboratory at Princeton University. AP is supported by a grant from the Simons Foundation (MP-SCMPS-00001470). AP additionally acknowledges support by NASA grant 80NSSC22K1054. ML was supported by the NASA Hubble Fellowship Program award HST-HF2-51537.002-A and NASA Astrophysics Theory Program award 80NSSC22K0817. This research benefited from interactions at the Kavli Institute for Theoretical Physics, supported by NSF PHY-2309135.
@@ Appendix A Numerical convergence
Figure 13: Comparison of the time evolution of the accretion rate $\dot{M}$ (a), magnetic flux at the horizon $\Phi_{\rm BH}$ (b), normalized dimensionless magnetic flux at the horizon $\phi_{\rm BH}$ (c), and efficiency $\eta$ (d) for the same accretion problem at different resolutions, represented by different colors. Dotted lines demonstrate simulation at a “medium” resolution performed by inclining the system by $90\deg$.
We perform the main simulation, $r_{b}=500r_{g}$, $\beta_{0}=10^{3}$, and $a=0.95$ at three different resolutions: small, medium (shown in the main text), and large. These correspond to grid sizes $384\times 144\times 256$, $812\times 384\times 512$, and $1536\times 768\times 1024$ in $N_{r}\times N_{\theta}\times N_{\phi}$. The time evolution of $\dot{M}$, $\Phi_{\rm BH}$, $\phi_{\rm BH}$, and $\eta$ are shown in Figure 13(a-d). Both small and medium resolutions show convergence with time. Due to high computational cost, we run the large resolution simulation to the point where $\phi_{\rm BH}$ falls, converging with other resolutions. Another uncertainty could arise due to the spherical grid’s polar singularity. We incline the BH spin by $90\deg$ at medium resolution to test this. Therefore, the jet launches perpendicular to the polar singularity. We see the convergence of main parameters relative to the grid inclination (Fig. 13a-d, dashed and solid purple lines).
@@ Appendix B Tetrad transformation
We utilize widely-used notations in a Boyer-Lindquist coordinates $t,r,\theta,\phi$ ($M\equiv c\equiv 1$):
@@
Covariant basis vectors of local non-rotating frame (LNRF) in BL coordinates (Bardeen et al., 1972):
@@
Transforming these basis vectors to Kerr-Schild (KS) coordinates $\tilde{\mu}$ ($r$ and $\theta$ do not change from BL to KS):
@@
Finally, 3-velocity in tetrad basis of BL coordinates, transformed from 4-velocity $u^{\tilde{\mu}}$ in KS coordinates:
@@
Keplerian 4-velocity in BL coordinates is $u_{\rm K}^{\mu}=u^{t}(\delta_{t}^{\mu}+\Omega_{\rm K}\delta_{\phi}^{\mu})$, where $\Omega_{\rm K}=(a+r^{3/2})^{-1}$. Using the tetrad basis for transforming the Keplerian contravariant velocity in BL coordinates to LNRF:

### v2

@@ 4 Conclusion and discussion
We have restricted our study to an initial magnetic field in the laboratory frame that is strictly poloidal and aligned with the spin axis of the BH. However, simulations with a Bondi radius of $200r_{g}$ and zero angular momentum (Ressler et al., 2021) have shown that tilting the magnetic field affects the resulting jet power. It remains unclear whether this carries over to the larger Bondi radius regime of most astrophysical interest. Therefore, in future studies, it would be interesting to explore the impact of tilting the magnetic field, as well as gas rotation, relative to the spin axis of the BH. These model problems could help in understanding the connection between the finite circularization radius problem studied here and simulations that begin with an equilibrium torus initial condition, in which tilting the magnetic field and plasma rotation axis relative to the BH spin axis can dramatically change the dynamics (e.g., Liska et al., 2018, 2019). Inflowing gas with a finite circularization radius likely serves as a better initial condition for the accretion of hot gas in the centers of galaxies than an equilibrium torus (e.g., Cuadra et al., 2006; Guo et al., 2024).
@@
The results presented here on the interplay between accretion and jets in GRMHD simulations of spherical accretion build on those of Ressler et al. (2021), Kwan et al. (2023), and Lalakos et al. (2024); we have extended these author’s results by varying the black hole spin, the initial gas angular momentum, the Bondi radius, and the initial plasma-$\beta$. For zero gas angular momentum, we find a similar dynamical evolution to that reported by Lalakos et al. (2024) – the initial emergence of a powerful jet with $\eta>100\%$, followed by its eventual disappearance as the normalized magnetic flux at the horizon ($\phi_{\rm BH}$) saturates below the typical MAD value. However, we attribute this behavior to the turbulence driven by eruption episodes, rather than the jet’s current-driven kink instability, as proposed by Lalakos et al. (2024). This interpretation is supported by our investigation of the $a=0$ case, where no jets are produced. Even in the absence of jets, the coherence of the accreted magnetic field is disturbed, leading to less efficient magnetic flux accumulation. In addition, if the transitions from $\eta>100\%$ to a low $\eta$ state were driven by jet feedback on large scales, we would expect the transition to take longer for a larger Bondi radius and for the accumulated magnetic flux to plausibly depend on the Bondi radius. Neither of these is true in our simulations. For example, our results show that the duration of the high-$\eta$ jet phase depends weakly on the Bondi radius, as seen in the initial spike in $\eta$ in Fig. 9. This is consistent with the disruption of magnetic flux accumulation originating in feedback from magnetic eruptions close to the black hole, rather than from larger radii. Once the strong jet is disrupted, however, the remaining weaker jets are indeed more prone to kink instability, which generates weak jets that dissipate their energy over a large solid angle (Fig. 3).
@@
One possible application of our findings is to accretion onto Sgr A*. In their simulations of accretion onto Sgr A* from the surrounding stellar winds, Ressler et al. (2023) found periods of time with large dimensionless magnetic flux and strong jets, and periods of time with weak jets. These may be analogous to the high- and low-angular-momentum states found here. More broadly, our results suggest that in the low-angular-momentum state explored in this work, observations should indicate the presence of strong magnetic fields even in the absence of a powerful jet. This is potentially consistent with the EHT and GRAVITY evidence for strong magnetic fields in Sgr A* (GRAVITY Collaboration et al., 2018; Gravity Collaboration et al., 2023; Event Horizon Telescope Collaboration et al., 2024), but the lack of definitive evidence for a powerful jet. As a check on this interpretation, we performed ray-tracing calculations of the synchrotron emission in the strong and weak jet states found in this work (as described in § 3.1). We find that the linear and circular polarization levels are indeed similar in both cases. Low angular momentum in the inflowing gas could also help explain the surprising lack of azimuthal asymmetry in the EHT image of Sgr A* (Cole Faggert et al., 2024).
A.G. and A.P. thank Misha Medvedev for fruitful discussions; we also thank Sasha Tchekhovskoy for useful conversations. The simulations presented in this work were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center and Visualization Laboratory at Princeton University. A.P. is supported by a grant from the Simons Foundation (MP-SCMPS-00001470). A.P. also acknowledges support from NASA grant 80NSSC22K1054, NSF grants PHY-2231698 and PHY-2206610, by an Alfred P. Sloan Research Fellowship and a Packard Foundation Fellowship in Science and Engineering. M.L. was supported by the NASA Hubble Fellowship Program award HST-HF2-51537.002-A and NASA Astrophysics Theory Program award 80NSSC22K0817. This research benefited from interactions at the Kavli Institute for Theoretical Physics, supported by NSF PHY-2309135.
@@ Appendix A Numerical convergence
Figure 13: Comparison of the time evolution of the accretion rate $\dot{M}$ (a), magnetic flux at the horizon $\Phi_{\rm BH}$ (b), normalized dimensionless magnetic flux at the horizon $\phi_{\rm BH}$ (c), and efficiency $\eta$ (d) for the same accretion problem at different resolutions, represented by different colors. The dotted lines demonstrate simulation at a “medium” resolution performed by inclining the system by $90\deg$.
We perform the main simulation, $r_{b}=500r_{g}$, $\beta_{0}=10^{3}$, and $a=0.95$ at three different resolutions: small, medium (shown in the main text), and large. These correspond to grid sizes $384\times 144\times 256$, $812\times 384\times 512$, and $1536\times 768\times 1024$ in $N_{r}\times N_{\theta}\times N_{\phi}$. The time evolution of $\dot{M}$, $\Phi_{\rm BH}$, $\phi_{\rm BH}$, and $\eta$ are shown in Figures 13((a)-(d)). Both small and medium resolutions show convergence with time. Due to high computational cost, we run the large resolution simulation to the point where $\phi_{\rm BH}$ falls, converging with other resolutions. Another uncertainty could arise due to the spherical grid’s polar singularity. We incline the BH spin by $90\deg$ at medium resolution to test this. Therefore, the jet launches perpendicular to the polar singularity. We see the convergence of main parameters relative to the grid inclination (Figures 13(a)-(d), dashed and solid purple lines).
@@ Appendix B Tetrad transformation
We utilize widely used notations in a Boyer-Lindquist (BL) coordinates $t,r,\theta,\phi$ ($M\equiv c\equiv 1$):
@@
Covariant basis vectors of the LNRF in BL coordinates (Bardeen et al., 1972):
@@
Transforming these basis vectors to KS coordinates $\tilde{\mu}$ ($r$ and $\theta$ do not change from BL to KS):
@@
Finally, three-velocity in tetrad basis of BL coordinates, transformed from four-velocity $u^{\tilde{\mu}}$ in KS coordinates:
@@
The Keplerian four-velocity in BL coordinates is $u_{\rm K}^{\mu}=u^{t}(\delta_{t}^{\mu}+\Omega_{\rm K}\delta_{\phi}^{\mu})$, where $\Omega_{\rm K}=(a+r^{3/2})^{-1}$. Using the tetrad basis for transforming the Keplerian contravariant velocity in BL coordinates to LNRF: